# Recovering contact forms from boundary data

### v5

@@ 4. Contact Holography: recovering contact geometry from the boundary data
Remark 4.1. Proposition 4.1 suggests that should be a geometric connection between cohomology classes with the values in the groups $H^{\ast+1}(X;\pi_{\ast}(SO(2n)/U(n))$ that obstruct the existence of complex (symplectic) structure on $\xi(v)$ and the geometry of the loci in $\partial_{1}^{+}X(v)$, where $(d\beta)^{n}|_{\partial_{1}^{+}X(v)}$ changes sign. Note that this change in sign can be expressed as the global property $\int_{{\partial_{1}^{+}X(v)}}|(d\beta)^{n}|>\int_{{\partial_{1}^{+}X(v)}}(d\beta)^{n}$. $\diamondsuit$
As Corollary 4.5 testifies, not any traversing vector field $v$ is the Reeb field for some contact form $\beta$! Some obstructions for finding such contact forms $\beta$ are of the homotopy-theoretical nature, others (see Proposition 4.2) are more obscure.
At the first glance, it seems probable that any traversing vector field $v$ on a $(2n+1)$dimensional manifold $X$ is a Reeb vector field for a contact form $\beta$, provided that a transversal to $v$ $(2n)$-distribution $\xi$ admits a complex/symplectic structure. However, this conjecture is false, as the next simple example testifies.
[…]
Let us generalize a bit Example 4.1.
@@
A $2k$-shell is the product $S\times[0,1]$, where $S$ is a connected oriented closed smooth manifold of dimension $2k$, $k\leq n$.
Let $v$ be a smooth non-vanishing vector field on a $(2n+1)$-dimensional compact manifold $X$ and let $\xi$ be a $(2n)$-distribution, transversal to $v$ and admitting a complex/symplectic structure.
We say that $(X,v,\xi)$ contains a contact $(2k+1)$-shell if there is a regular embedding $\Psi:S\times[0,1]\to X$ that maps each oriented segment $s\times[0,1]$, $s\in S$, to an oriented segment of a $v$-trajectory and such that a normal $(2n-2k)$-bundle $\nu(\Psi)$ of $\Psi(S\times[0,1])$ in $X$ is a complex/symplectic subbundle of the $2n$-bundle $\xi|_{\Psi(S\times[0,1])}$. $\diamondsuit$
Equivalently, $(X,v,\xi)$ contains a contact $(2k+1)$-shell $\Psi:S\times[0,1]\to X$ if and only if there is a regular embedding $\psi:S\to X$ such that there exists a symplectic $(2n-2k)$-subbundle $\mu$ of $\xi|_{\psi(S)}$ whose fibers are transversal to $\psi(S)$; i.e., $\mathbb{R}_{v}\oplus\mu$ is a normal bundle of $\psi(S)$ in $X$, where $\mathbb{R}_{v}$ stands for the line bundle spanned by $v$.
@@
Let $\beta$ be a contact form whose Reeb field is $v_{\beta}$. Let $\xi_{\beta}=\ker(\beta)$. Assume that $(X,v_{\beta},\xi_{\beta})$ contains a contact $(2k)$-shell $\Psi(S\times[0,1])$. Thanks to Definition 4.3, the normal bundle $\nu(\Psi)$ is a complex/symplectic subbundle of $\xi_{\beta}|_{\Psi(S\times[0,1])}$.
Consider the form $\Theta^{\bullet}=\Psi^{\ast}(\beta\wedge(d\beta)^{k})$ on $S\times[0,1]$. Since $\beta\wedge(d\beta)^{n}$ is a volume form and $(d\beta)^{n-k}|_{\nu(\Psi)}$ is the volume form by Definition 4.3, the form $\Theta^{\bullet}$ must be a volume form on $S\times[0,1]$. Thus, $\beta^{\bullet}=_{\mathsf{def}}\Psi^{\ast}(\beta)$ is a contact form on $S\times[0,1]$ and the transfer $v^{\dagger}$ of $v_{\beta}$ under $\Psi$ is the Reeb vector field for $\beta^{\bullet}$. Moreover, $\partial_{1}^{+}(S\times[0,1])(v^{\dagger})=S\times\{0\}$.
@@
Let $\beta$ be a contact form on an oriented smooth $(2n+1)$-manifold $X$. Let the Reeb vector field $v_{\beta}$ of $\beta$ be transversal to a smooth connected oriented hypersurface $Y\subset X$, an open $(2n)$-manifold.
@@
Let $U$ be a smooth regular neighborhood of $Y$ in $X$. Consider the $(-v_{\beta})$ induced projection $Q^{\beta}:U\to Y$. We denote its differential by $Q^{\beta}_{\ast}$. Since $v_{\beta}$ is transversal to $Y$, the map $Q^{\beta}_{\ast}|:\xi_{\beta}|_{Y}\to T(Y)$ is a symplectomorphism with respect to forms $d\beta|_{\xi_{\beta}}$ and $d\beta|_{Y}$. If the normal bundle $\nu(S)$ of $S$ in $Y$ is a symplectic subbundle of $TY$, then $(Q^{\beta}_{\ast})^{-1}\big{(}\nu(S)\big{)}\cap\xi_{\beta}$ is a symplectic subbundle of $\xi_{\beta}|_{(Q^{\beta})^{-1}(S)}$. It is on the level of definitions to check that $(Q^{\beta})^{-1}(S)$ is a contact shell in $U\subset X$. Now the corollary follows from Proposition 4.2. ∎
@@
Let $\mathcal{V}_{\{\neq 0,\;\xi\text{ is symp.}\}}(X)$ be the space of smooth non-vanishing vector fields $v$ on $X$ that admit a transversal to them symplectic $(2n)$-distribution $\xi\subset TX$, where $\dim(X)=2n+1$. Let $k\in[1,n]$.
@@
Therefore, if $\mathcal{W}_{k}(X)\neq\emptyset$, then $\mathcal{V}_{\{\neq 0,\;\xi\text{ is symp.}\}}(X)$ contains an open nonempty set of pairs $(v,\xi)$, where $v$ is not the Reeb vector field for any contact form $\beta$ on $X$, subject to to $(d\beta)^{n}|_{\xi}>0$.
@@
Let $Z$ be a compact smooth $(2n+1)$-dimensional manifold and $f:Z\to\mathbb{R}$ a Morse function. Let $\tilde{v}$ be a $f$-gradient-like vector field on $Z$. We denote by $X$ the complement in $Z$ to a disjoint union of small open $(2n+1)$-balls $\{B_{z}\}$ that are centered on the critical points $\{z\}$ of $f$ and are convex in the local euclidean metric $g_{z}$, adjusted to the local Morse coordinates at $z$. A similar construction makes sense for any closed $1$-form $\alpha$ on $Z$ with Morse type singularities. In their vicinity, $\alpha=df$, where $f$ has the standard Morse form in the Morse coordinates at each singularity. Recall that a vector field $v$ is of the gradient type for $\alpha$ on $X=Z\setminus\bigcup_{z\in\mathsf{sing}(\alpha)}B_{z}$ if $\alpha(v)>0$ on $X$.
@@
If $Z$ is a closed manifold, then the gradient-like vector field $v$ on the manifold $X=Z\setminus\bigcup_{z\in\mathsf{crit}(f)}B_{z}$ is not a Reeb field of any contact form.
@@
For any $f$-critical point $z$, consider the sphere $S^{2n}_{z}=\partial B_{z}$. Let $a\in H^{2n}(S^{2n}_{z};\mathbb{Z})$ be a generator. Let $\tilde{K}(S^{2n})$ denote the reduced complex $K$-theory of the sphere.
For each complex vector bundle $\eta$ over $S^{2n}$, the $n$-th Chern class $c_{n}(\eta)$ is a multiple of $(n-1)!\,a$. Moreover, for each $m$ such that $m\equiv 0\mod(n-1)!$, there is a unique $[\eta]\in\tilde{K}(S^{2n})$ such that $c_{n}(\eta)=ma$ (see [Huse], 9.8 Corollary).
[…]
@@
Unfortunately, this arithmetic condition can be satisfied only for $n\leq 3$. Therefore, no gradient field $v$ of a Morse function $f:Z\to\mathbb{R}$ can produce a contact form $\beta$ on $X=Z\setminus(\bigcup_{z\in\mathsf{crit}(f)}\;D_{z})$ such that $v_{\beta}=v|_{X}$ in all dimensions $\dim X\geq 9$.
@@
Therefore, $v$ is concave in relation to $\partial_{1}X$, and Proposition 4.1 is applicable. Hence, by the argument above, for any $1$-form $\beta$ such that $\beta(v)=1$ and $d\beta(v)=0$, the form $(d\beta)^{n}\big{|}_{\mathsf{int}(\partial_{1}^{+}X(v))}$ must flip sign, provided $\dim X\geq 9$. ∎
@@
Let $\tilde{v}$ be a vector field on an odd-dimensional manifold $Y$ such that the zeros of $\tilde{v}$ form a finite set $\mathcal{Z}_{\tilde{v}}$. Let $U(\mathcal{Z}_{\tilde{v}})$ be an open regular neighborhood of $\mathcal{Z}_{\tilde{v}}$ in $Y$. Put $X=Z\setminus U(\mathcal{Z}_{\tilde{v}})$ and let $v=\tilde{v}|_{X}$.
@@
[…]
@@
Then there is a smooth isotopy of $\{\phi_{t}:X\to X\}_{t\in[0,1]}$ such that the Reeb vector field $v_{\beta_{1}}$ of the contact form $\beta_{1}=\phi_{1}^{\ast}(\beta)$ is traversing on $\phi_{1}(X)\subset X$. $\diamondsuit$
Since a connected manifold with boundary admits a handle decomposition without handles of the top dimension $12n+$ 1, there is no homotopy-theoretical obstruction for extending any almost contact structure (see Definition 4.1) from the base of each handle to its interior (so that the preconditions for applying the $h$-principle are valid). Therefore, by Gromov’s Theorem [Gr1] (see also its enforcement, the fundamental Theorem 1.1 of Borman, Eliashberg, and Murphy [BEM]), the cooriented contact structure $\xi$ on $X$ exists. However, we do not know whether the appropriate transversal to $v_{\beta}$ and $\hat{v}_{\beta}$-invariant foliation $\mathcal{G}$ exists.
For a traversing vector field $v$ on $X$, every $v$-trajectory $\gamma\subset X$ has a neighborhood $\hat{U}(\gamma)\subset\hat{X}$ and a $t$-parametric family of diffeomorphisms $\{\phi^{t}:\hat{U}(\gamma)\to\hat{U}(\gamma)\}_{t\in[-\epsilon,\epsilon]}$, where $\epsilon>0$ depends on $\gamma$, such that $\phi^{0}=\mathsf{id}$, and $\frac{d}{dt}\phi^{t}(a)\big{|}_{t=0}=v(a)$ for all $a\in\hat{U}(\gamma)\cap X$.
@@
Let $v$ be a vector field on $X$. We say that a contact form $\beta$ is $v$-invariant if $(\phi^{t})^{\ast}(\beta)=\beta$ in $\hat{U}(\gamma)\cap X\cap(\phi^{t})(X)$ for all sufficiently small $t$. Equivalently, $\mathcal{L}_{v}\beta=0$.
@@
[…]
@@
Given a contact form $\beta$, if a smooth positive function $\lambda:X\to\mathbb{R}_{+}$ is such that $\mathcal{L}_{v_{\beta}}\lambda=0$, then the form $\beta_{1}=_{\mathsf{def}}\lambda^{-1}\beta$ is $\mathcal{RC}$-equivalent to the form $\beta$. Moreover, the Reeb vector field $v_{\beta_{1}}=\lambda v_{\beta}$. As a result, positive functions from the algebra $C^{\infty}(\mathcal{T}(v_{\beta}))$ of smooth functions on the trajectory space $\mathcal{T}(v_{\beta})$ act on $\beta$ by multiplication and transform it into contact forms $\mathcal{RC}$-equivalent to $\beta$. $\diamondsuit$
Remark 4.2.
Evidently, $\mathcal{RC}$-equivalent contact forms share the dynamics of their Reeb flows. $\diamondsuit$
@@
Let $\eta:X\to\mathbb{R}$ be a smooth function such that $\mathcal{L}_{v_{\beta}}\eta=d\eta(v_{\beta})=0$. Consider the $1$-form $\beta_{1}=\beta+d\eta$. Then $\beta_{1}$ is $\mathcal{R}$-equivalent to $\beta$.
@@
By definition, $d\eta(v_{\beta})=0,v_{\beta}\,\rfloor\,d\beta=0$, and for dimensional reasons $d\eta\wedge(d\beta)^{n}|_{\ker\beta}=0$. Thus, $d\eta\wedge(d\beta)^{n}=0$. Therefore, $\beta_{1}\wedge(d\beta_{1})^{n}=(\beta+d\eta)\wedge(d\beta)^{n}=\beta\wedge(d\beta)^{n}>0$. Also, $\beta_{1}(v_{\beta})=\beta(v_{\beta})+d\eta(v_{\beta})=1$ and, since $d\beta_{1}=d\beta$, we get $v_{\beta}\,\rfloor\,d\beta_{1}=0$. Hence, $v_{\beta_{1}}=v_{\beta}$. ∎

### v6

@@ 4. Contact Holography: recovering contact geometry from the boundary data
Remark 4.1. Proposition 4.1 suggests that should be a geometric connection between cohomology classes with the values in the groups $H^{\ast+1}(X;\pi_{\ast}(\mathsf{SO}(2n)/\mathsf{U}(n))$ that obstruct the existence of complex (symplectic) structure on $\xi(v)$ and the geometry of the loci in $\partial_{1}^{+}X(v)$, where $(d\beta)^{n}|_{\partial_{1}^{+}X(v)}$ changes sign. Note that this change in sign can be expressed as the global property $\int_{{\partial_{1}^{+}X(v)}}|(d\beta)^{n}|>\int_{{\partial_{1}^{+}X(v)}}(d\beta)^{n}$. $\diamondsuit$
As Corollary 4.5 testifies, not any traversing vector field $v$ is the Reeb field for some contact form $\beta$! Some obstructions for finding such contact forms $\beta$ are of the homotopy-theoretical nature, others (see Proposition 4.2 below) are more obscure.
At the first glance, it seems probable that any traversing vector field $v$ on a $(2n+1)$ -dimensional manifold $X$ is a Reeb vector field for a contact form $\beta$, provided that a transversal to $v$ $(2n)$-distribution $\xi$ admits a complex/symplectic structure. However, this conjecture is false, as the next simple example testifies.
Example 4.1. Consider the $3$-fold $X=S\times[0,1]$ where $S$ is an oriented closed surface. Let $v$ be a constant vector on $X$, tangent to the fibers of the projection $S\times[0,1]\to S$. Then Proposition 4.1 tells us that such a $v$ is not a Reeb vector field of any contact form $\beta$ on $X$. Indeed, any transversal to $v$ plane field $\xi$ is orientable and thus admits a complex/symplectic structure. However, $d\beta$ cannot be positive on $\partial^{+}_{1}X(v)=S\times\{0\}$ since $\int_{S}d\beta=0$ by Stokes’ Theorem. $\diamondsuit$
Let us generalize Example 4.1.
@@
A $2k$-shell is the product $S\times[0,1]$, where $S$ is a connected oriented closed smooth manifold of a dimension $2k$, $k\leq n$.
Let $v$ be a smooth non-vanishing vector field on a $(2n+1)$-dimensional compact manifold $X$, and let $\xi$ be a $(2n)$-distribution, transversal to $v$ and admitting a complex/symplectic structure.
We say that $(X,v,\xi)$ contains a contact $(2k+1)$-shell if there is a regular embedding $\Psi:S\times[0,1]\to X$ that maps each oriented segment $s\times[0,1]$, $s\in S$, to an oriented segment of a $v$-trajectory and such that a normal (2n2k)(2n-2k)-bundle $\nu(\Psi)$ of $\Psi(S\times[0,1])$ in $X$ is a complex/symplectic subbundle of the $2n$-bundle $\xi|_{\Psi(S\times[0,1])}$ (viewed a $n$-dimensional complex bundle). $\diamondsuit$
Equivalently, $(X,v,\xi)$ contains a contact $(2k+1)$-shell $\Psi:S\times[0,1]\to X$ if and only if there is a regular embedding $\psi:S\to X$ such that there exists a symplectic (2n2k)(2n-2k)-subbundle $\mu$ of $\xi|_{\psi(S)}$ whose fibers are transversal to $\psi(S)$; i.e., $\mathbb{R}_{v}\oplus\mu$ is a normal bundle of $\psi(S)$ in $X$, where $\mathbb{R}_{v}$ stands for the line bundle spanned by $v$.
@@
Let $\beta$ be a contact form whose Reeb field is $v_{\beta}$. Let $\xi_{\beta}=\ker(\beta)$. Assume that $(X,v_{\beta},\xi_{\beta})$ contains a contact $(2k+1)$-shell $\Psi(S\times[0,1])$. Thanks to Definition 4.3, the normal bundle $\nu(\Psi)$ of $\Psi(S\times[0,1])$ is a complex/symplectic subbundle of $\xi_{\beta}|_{\Psi(S\times[0,1])}$.
Consider the form $\Theta^{\bullet}=\Psi^{\ast}(\beta\wedge(d\beta)^{k})$ on $S\times[0,1]$. Since $\beta\wedge(d\beta)^{n}$ is a volume form and $(d\beta)^{n-k}|_{\nu(\Psi)}$ is the volume form by Definition 4.3, the form $\Theta^{\bullet}$ must be a volume form on $S\times[0,1]$. Thus, $\beta^{\bullet}=_{\mathsf{def}}\Psi^{\ast}(\beta)$ is a contact form on $S\times[0,1]$ and the transfer $v^{\dagger}$ of $v_{\beta}$ under $\Psi$ is the Reeb vector field of $\beta^{\bullet}$. Moreover, $\partial_{1}^{+}(S\times[0,1])(v^{\dagger})=S\times\{0\}$.
@@
Let $\beta$ be a contact form on a smooth $(2n+1)$-manifold $X$. Let the Reeb vector field $v_{\beta}$ be transversal to a smooth connected oriented hypersurface $Y\subset X$, an open $(2n)$-manifold.
@@
Let $U$ be a smooth regular neighborhood of $Y$ in $X$. Consider the $(-v_{\beta})$ induced projection $Q^{\beta}:U\to Y$. We denote its differential by $Q^{\beta}_{\ast}$. Since $v_{\beta}$ is transversal to $Y$, the isomorphism $Q^{\beta}_{\ast}|:\xi_{\beta}|_{Y}\to T(Y)$ is a symplectomorphism with respect to the forms $d\beta|_{\xi_{\beta}}$ and $d\beta|_{Y}$. If the normal bundle $\nu(S)$ of $S$ in $Y$ is a symplectic subbundle of $TY$, then $(Q^{\beta}_{\ast})^{-1}\big(\nu(S)\big)\cap\xi_{\beta}$ is a symplectic subbundle of $\xi_{\beta}|_{(Q^{\beta})^{-1}(S)}$. It is on the level of definitions to check that $(Q^{\beta})^{-1}(S)$ is a contact shell in $U\subset X$. Now the corollary follows from Proposition 4.2. ∎
@@
Let $\mathcal{V}_{\{\neq 0,\;\xi\text{ is symp.}\}}(X)$ be the space of all smooth non-vanishing vector fields $v$ on a $(2n+1)$-dimensional manifold $X$ that admit a transversal to $v$ symplectic $(2n)$-distribution $\xi\subset TX$. Let $k\in[1,n]$.
@@
Therefore, if $\mathcal{W}_{k}(X)\neq\emptyset$, then $\mathcal{V}_{\{\neq 0,\;\xi\text{ is symp.}\}}(X)$ contains an open nonempty set of pairs $(v,\xi)$ such that $v$ is not the Reeb vector field for any contact form $\beta$ on $X$, subject to to $(d\beta)^{n}|_{\xi}>0$.
@@
Let $Z$ be a compact smooth $(2n+1)$-dimensional manifold and $f:Z\to\mathbb{R}$ a Morse function. Let $\tilde{v}$ be a $f$-gradient-like vector field on $Z$. We denote by $X$ the complement in $Z$ to a disjoint union of small open $(2n+1)$-balls $\{B_{z}\}$ that are centered on the critical points $\{z\}$ of $f$ and are convex in the local Euclidean metric $g_{z}$, adjusted to the local Morse coordinates at $z$. A similar construction makes sense for any closed $1$-form $\alpha$ on $Z$ with the Morse type singularities. In their vicinity, $\alpha=df$, where $f$ has the standard Morse form in the Morse coordinates at each singularity. Recall that a vector field $v$ is of the gradient type for $\alpha$ on $X=Z\setminus\bigcup_{z\in\mathsf{sing}(\alpha)}B_{z}$ if $\alpha(v)>0$ on $X$.
@@
If $Z$ is a closed odd-dimensional manifold, then the gradient-like vector field $v$ on the manifold $X=Z\setminus\bigcup_{z\in\mathsf{crit}(f)}B_{z}$ is not a Reeb field of any contact form.
@@
For any $f$-critical point $z$, consider the sphere $S^{2n}_{z}=\partial B_{z}$. Let $a\in H^{2n}(S^{2n}_{z};\mathbb{Z})$ be a generator. Let $\tilde{K}(S^{2n})$ denote the reduced complex $K$-theory of the sphere $S^{2n}$.
For each complex vector bundle $\eta$ over $S^{2n}$, the $n$-th Chern class $c_{n}(\eta)$ is a multiple of (n1)!a(n-1)!\,a. Moreover, for each $m$ such that $m\equiv 0\mod(n-1)!$, there is a unique $\eta]\in\tilde{K}(S^{2n})$ such that $c_{n}(\eta)=ma$ (see [Huse], 9.8 Corollary).
On the other hand, for any Morse singularity $z$, the Euler class of a $(2n)$-dimensional distribution $\xi(v)|_{S^{2n}_{z}}$, tangent to the hypersurfaces of $f$-constant level, can be computed by the formula $\chi(\xi(v))=(-1)^{\mathsf{ind}_{z}(f)}\,2$, where $\mathsf{ind}_{z}(f)$ is the Morse index of $z$. This can be seen by applying the Gaussian map to $v$ along $S^{2n}_{z}$, while considering the orthogonal to $v$ hyperplane field $v^{\perp}\approx\xi(v)|_{S^{2n}_{z}}$. Therefore, if the real $2n$-bundle $\xi(v)$ has a complex structure, then for each $f$-critical point $z$, we must have $(-1)^{\mathsf{ind}_{z}(f)}\,2\equiv 0\mod(n-1)!$. As a result, a necessary condition for this construction to lead to a contact structure on the $(2n+1)$-dimensional $X$ is
@@
Unfortunately, this arithmetic condition can be satisfied only for $n\leq 3$. Therefore, no gradient field $v$ of a Morse function $f:Z\to\mathbb{R}$ is amenable to any contact form $\beta$ on $X=Z\setminus(\bigcup_{z\in\mathsf{crit}(f)}\;D_{z})$ such that $v_{\beta}=v|_{X}$ in all dimensions $\dim X\geq 9$.
@@
Therefore, $v$ is concave in relation to $\partial_{1}X$ (i.e., $\partial_{2}^{-}X(v)=\emptyset$), and Proposition 4.1 is applicable. Hence, by the argument above, for any $1$-form $\beta$ such that $\beta(v)=1$ and $d\beta(v)=0$, the form $(d\beta)^{n}\big|_{\mathsf{int}(\partial_{1}^{+}X(v))}$ must flip sign, provided $\dim X\geq 9$.
@@
[…]
@@
$(-1)^{\mathsf{ind}_{z}(v)}\,2\equiv\;0\mod(n-1)!\text{ for each }z\in\mathcal{Z}_{\tilde{v}}$,
@@
Then there is a smooth isotopy of $\{\phi_{t}:X\to X\}_{t\in[0,1]}$ such that the Reeb vector field $v_{\beta_{1}}$ of the contact form $\beta_{1}=\phi_{1}^{\ast}(\beta)$ is traversing (see Definition 2.1) on $\phi_{1}(X)\subset X$. $\diamondsuit$
Remark 4.1.
Since a connected manifold with boundary admits a handle decomposition without handles of the top dimension $2n+1$, there is no homotopy-theoretical obstruction for extending any almost contact structure (see Definition 4.1) from the base of each handle to its interior (so that the preconditions for applying the $h$-principle are valid). Therefore, by Gromov’s Theorem [Gr1] (see also its enforcement, the fundamental Theorem 1.1 of Borman, Eliashberg, and Murphy [BEM]), the cooriented contact structure $\xi$ on $X$ exists. However, we do not know whether the appropriate transversal to $v_{\beta}$ and $\hat{v}_{\beta}$-invariant foliation $\mathcal{G}$ exists. $\diamondsuit$
For a traversing vector field $v$ on $X$, every $v$-trajectory $\gamma\subset X$ has a neighborhood $\hat{U}(\gamma)\subset\hat{X}$ and a $t$-parametric family of diffeomorphisms $\{\phi^{t}:\hat{U}(\gamma)\to\hat{U}(\gamma)\}_{t\in[-\epsilon,\epsilon]}$ (where $\epsilon>0$ depends on $\gamma$) such that $\phi^{0}=\mathsf{id}$, and $\frac{d}{dt}\phi^{t}(a)\big|_{t=0}=v(a)$ for all $a\in\hat{U}(\gamma)\cap X$.
@@
Let $v$ be a vector field on $X$. We say that a contact form $\beta$ is $v$-invariant if $(\phi^{t})^{\ast}(\beta)=\beta$ in $\hat{U}(\gamma)\cap X\cap(\phi^{t})(X)$ for all sufficiently small $t$. Equivalently, the directional derivative $\mathcal{L}_{v}\beta=0$.
@@
  • We call two contact forms $\beta_{1}$ and $\beta_{2}$ on $X$ Reeb-equivalent ($\mathcal{R}$-equivalent for short), if they share the same Reeb vector field $v$. Similarly, we call contact structures $\xi_{1}$ and $\xi_{2}$ $\mathcal{R}$-equivalent if they are are generated by $\mathcal{R}$-equivalent contact forms.
@@
Evidently, $\mathcal{RC}$-equivalent contact forms $\beta_{1}$ and $\beta_{2}$ share the dynamics of their Reeb flows. In particular, if $v_{\beta_{1}}$ is traversing, or boundary generic, or admits a Lyapunov $1$-form/function, then so is/does $v_{\beta_{2}}$. $\diamondsuit$
@@
Given a contact form $\beta$ on $X$ and a smooth positive function $\lambda:X\to\mathbb{R}_{+}$ such that $\mathcal{L}_{v_{\beta}}\lambda=0$, then the form $\beta_{1}=_{\mathsf{def}}\lambda\beta$ is $\mathcal{RC}$-equivalent to the form $\beta$. Moreover, the Reeb vector field $v_{\beta_{1}}=\lambda^{-1}v_{\beta}$. $\diamondsuit$
@@
Evidently, $(\lambda^{-1}v_{\beta})\rfloor(\lambda\beta)=1$. Using that $d(\lambda\beta)=d\lambda\wedge\beta+\lambda d\beta$, we get
$(\lambda^{-1}v_{\beta})\rfloor(d\lambda\wedge\beta+\lambda d\beta)=\lambda^{-1}d\lambda(v_{\beta})\beta+v_{\beta}\rfloor d\beta=0$
since $d\lambda(v_{\beta})=\mathcal{L}_{v_{\beta}}\lambda=0$ by the choice of $\lambda$ and $v_{\beta}\rfloor d\beta=0$. Therefore, $\lambda^{-1}v_{\beta}$ is the Reeb field for $\lambda\beta$.
As a result, for a traversing Reeb field $v_{\beta}$, positive functions from the algebra $C^{\infty}(\mathcal{T}(v_{\beta}))$ of smooth functions on the trajectory space $\mathcal{T}(v_{\beta})$ act on $\beta$ by multiplication and transform it into contact forms $\mathcal{RC}$-equivalent to $\beta$. ∎
Lemma 4.6.
Let $\beta$ be a contact form on a smooth manifold $X$ and $v_{\beta}$ its Reeb vector field. Consider a $1$-form $\beta_{1}=\beta+d\eta$, where the smooth function $\eta:X\to\mathbb{R}$ is such that $d\eta(v_{\beta})>-1$.222i.e., $\eta$ is not “negatively very steep” along the Reeb flow trajectories. Such functions $\eta$ form an open convex set in $C^{\infty}(X)$.
Then the form $\beta_{1}$ is a contact form on $X$. Moreover, $\beta_{1}$ is $\mathcal{RC}$-equivalent to $\beta$.
Proof.
Since $\beta_{1}=\beta+d\eta$, we get
$\beta_{1}\wedge(d\beta_{1})^{n}=\beta_{1}\wedge(d\beta)^{n}=\beta\wedge(d\beta)^{n}+d\eta\wedge(d\beta)^{n}.$
We claim that the $(2n+1)$-form $\beta_{1}\wedge(d\beta_{1})^{n}$ is positive on $X$, provided $d\eta(v_{\beta})>-1$. Indeed, the $(2n+1)$-form $d\eta\wedge(d\beta)^{n}$ is proportional to $(2n+1)$-form $\beta\wedge(d\beta)^{n}$ with some functional coefficient $h:X\to\mathbb{R}$. Thus $\beta_{1}\wedge(d\beta_{1})^{n}>0$ is equivalent to the inequality $1+h>0$ on $X$.
To detect $h$, it suffices to compare the contractions of the two forms, $d\eta\wedge(d\beta)^{n}=h(\beta\wedge(d\beta)^{n})$ and $\beta\wedge(d\beta)^{n}$, with a non-vanishing Reeb vector field $v_{\beta}$. Recall that $v_{\beta}\rfloor\beta=1$ and $v_{\beta}\rfloor d\beta\equiv 0$. Using these properties of $v_{\beta}$, we get
$v_{\beta}\rfloor(\beta\wedge(d\beta)^{n})=(d\beta)^{n},\text{ while }v_{\beta}\rfloor(d\eta\wedge(d\beta)^{n})=d\eta(v_{\beta})(d\beta)^{n}=h\;(d\beta)^{n}.$
Thus, when $d\eta(v_{\beta})>-1$, $v_{\beta}\rfloor(\beta_{1}\wedge(d\beta_{1})^{n})$ is positively proportional to $(d\beta)^{n}$ with the functional coefficient $h$. Therefore, $\beta_{1}$ is a contact form, provided $d\eta(v_{\beta})>-1$.
Put $v_{\beta_{1}}=_{\mathsf{def}}(1+d\eta(v_{\beta}))^{-1}v_{\beta}$. Let us check that $v_{\beta_{1}}$ is the Reeb vector field for the form $\beta_{1}$. Indeed, $\beta_{1}(v_{\beta_{1}})=\beta_{1}((1+d\eta(v_{\beta}))^{-1}v_{\beta})=(\beta+d\eta)\big((1+d\eta(v_{\beta}))^{-1}v_{\beta}\big)=1$ since $\beta(v_{\beta})=1$, and $d\beta_{1}(v_{\beta_{1}})=d\beta((1+d\eta(v_{\beta}))^{-1}v_{\beta})=(1+d\eta(v_{\beta}))^{-1}d\beta(v_{\beta})\equiv 0.$ Thus, $\beta_{1}$ and $\beta$ are $\mathcal{RC}$-equivalent. As a result, if $v_{\beta}$ admits a Lyapunov function/form, then so does $v_{\beta_{1}}$. ∎
Combining Lemma 4.6 with Lemma 4.5 delivers the following corollary.
Corollary 4.7.
Given a contact form $\beta$ and smooth functions $\lambda$, $\eta$ such that $\mathcal{L}_{v_{\beta}}\lambda=0$ and $d\eta(v_{\beta})>-1$, the form $\beta_{1}=\lambda(\beta+d\eta)$ is a contact form. Moreover, $\beta_{1}$ and $\beta$ are $\mathcal{RC}$-equivalent. $\diamondsuit$